# Discovering and exploiting active sensing motifs for estimation

### v1

@@ 7. Reveal active sensing motifs by correlating state dynamics with observability
Adding a measurement of apparent airflow magnitude ($a$) made the offset turn sufficient to estimate wind direction and additionally allowed ground speed to be estimated (LABEL:fig:figure_2_cii). Furthermore, adding an optic flow magnitude measurement ($r$) made the acceleration/deceleration motifs viable for increasing the observability of wind direction and ground speed, and it rendered the altitude observable (LABEL:fig:figure_2_ciii).
Replacing the optic flow measurement with a direct ground speed measurement ($g$) rendered the wind direction and ground speed observable without any motif (even during straight flight), and thus active sensing would not be helpful for this sensor set (LABEL:fig:figure_2_civ). However, without optic flow this sensor set would not be sufficient to estimate altitude (LABEL:fig:figure_2_civ).
@@
A sensor set consisting of optic flow magnitude ($r$) and direction ($\beta$) and acceleration magnitude ($q$) and direction ($\nu$) was suitable for estimating ground speed and altitude during acceleration/deceleration and offset turns (LABEL:fig:figure_2_cv). However, wind direction was never observable and heading turns had minimal effect on the observability (LABEL:fig:figure_2_cv).
It is also important to consider the times at which measurements are collected when evaluating active sensing motifs. For some sets of measurements, information must be collected continuously throughout a motif to increase observability (LABEL:fig:figure_supp_timestep_a). However, other sets of measurements may allow information to be considered only before and after the motif, with a similar increase in observability to continuously collecting information (LABEL:fig:figure_supp_timestep_b). This is an especially important consideration when dealing with systems where sensor noise may be correlated with a motif. For instance, translational optic flow cues may be noisier or less reliable when turning due to the influence of rotational optic flow.
In summary, our approach reveals a diversity of active sensing motifs for a flying agent. We show how an agent would ideally employ distinct movement patterns based on the available measurements and the state variables being estimated. Our results are not particularly sensitive to the model type and generalize well when modeling more complex dynamics (Figure S2). Altogether, these motifs can be used to inform control/estimation strategies and algorithms for autonomous flight with limited sensors. BOUNDS can be applied in a similar manner to any arbitrary system to discover measurement- and movement-dependent sensing motifs.
@@
Figure 3: Observability-informed state estimation. a. Artificial neural network (ANN) state estimator for wind direction ($\mathcal{H}_{\zeta}$). Inputs are the yaw heading angle ($\psi$), apparent wind angle ($\gamma$), and course direction angle ($\beta$) in the time window $\omega$. b. A subset of simulated flying agent trajectories used for training and testing the ANN state estimator. $N=40,000$ trajectories total. c. The distribution (bottom) and cumulative distribution function (top) of wind direction observability ($\mathcal{F}^{-1}_{\zeta}$) across all training and testing trajectories. d. A comparison of the testing error variance for the wind direction state estimator ANN ($\mathcal{H}_{\zeta}$) when trained on different percentages of the training data—for data sorted by observability (sorted high-to-low, top) and randomly sorted (bottom). The y-axis indicates how much data was used in training, and the x-axis indicates what bin percentile was used to test the ANN (testing data always sorted by observability). e. Same as k. but showing the relative change in error variance between observability-sorted and randomly sorted training data. Red values (<< 0) indicate better performance when sorting by observability. f. The mean observability $\text{mean}(\mathcal{F}^{-1}_{\zeta})$ of each bin in the training dataset vs the ANN error variance $\text{Var}(\check{\zeta}-\zeta)$ for ANN’s trained on different percentages of the training dataset for observability-sorted (left) and randomly sorted (right) training data. The theoretical bound of the ANN performance is shown as a dashed line (bounded at 1 for circular data). g. (top) An example simulated trajectory of a flying agent in the presence of time-varying wind direction. Color along trajectory indicates observability. (bottom) A comparison of true wind direction ($\zeta$}), ANN wind direction estimate ($\check{\zeta}$), and the observability-filtered estimate ($\hat{\zeta}$). Measurement noise covariance was set to $\mathcal{R}=10^{-3}I_{p\times p}$.
@@
In this section, we present a framework for merging state and observability estimates for agents that rely on active sensing. We focus on four primary steps: 1) training a state estimator, 2) curating training data by observability, 3) training an observability estimator, and 4) observability-informed filtering fusing state and observability estimates. We show an example application of each step for wind direction estimation.
@@
$\mathcal{H}_{i}$ is a function that maps the time history of measurements $\mathbf{Y}_{\omega_{k}}$ (and optionally inputs $\mathbf{U}_{\omega_{k}}$ and measurement noise covariance $\mathcal{R}_{\omega_{k}}$) in the time window $\omega$ to the current state estimate $\check{x}_{i,k}$ at time step $k$. Notably, $\mathcal{H}_{i}$ can consider the entire time window $\omega$ of information at once, and is not sensitive to initialization, thus can potentially avoid the pitfalls associated with Kalman filtering. In principle, $\mathcal{H}_{i}$ could be any function, but we employ a feed-forward artificial neural network (ANN) for its generality, flexibility, and ability to operate without relying on a predefined dynamical model.
We designed a wind direction state estimator ANN using a limited set of angular measurements $\psi,\gamma,\beta]$ (LABEL:fig:figure_3_a, LABEL:fig:figure_supp_ANN_noise_aLABEL:fig:figure_supp_ANN_noise_b). Our observability analysis tells us that the agent must turn with a change in heading in order for the ANN to yield accurate estimates of wind direction (LABEL:fig:figure_2_ci).
@@
With a simulation model, or experimental testbed where ${x}_{i}$ can be measured directly, one should be able to generate a rich dataset to train and validate a state estimator $\mathcal{H}_{i}$. However, the distribution of observability within the training dataset is critical. For instance, attempting to train an estimator on primarily unobservable data—for instance, straight trajectories for our wind direction estimator (LABEL:fig:figure_2_ci)—will surely lead to poor results. Alternatively, one could leverage prior knowledge of observable motifs to generate high quality training data, or curate a subset of an existing dataset that has an appropriate distribution of observability.
We generated training data by randomly simulating a diversity of trajectories (LABEL:fig:figure_3_b, see Methods). We ensured that the trajectories densely spanned a large envelope of the state space and yielded a wide distribution of observability values (LABEL:fig:figure_3_c). We then separated the data into ten equally sized bins, sorted by the observability of wind direction (LABEL:fig:figure_3_c). This allowed us to train ANN’s on inputs with distinct distributions of observability. We specifically performed an analysis of the ANN wind direction estimate error variance $\text{Var}(\zeta-\check{\zeta})$ when trained on varying percentages of the training dataset—once when the training data was sorted by observability level, and once when the data was randomly sorted (LABEL:fig:figure_3_d, Figure S4).
When comparing the observability-sorted to randomly-sorted trained ANN’s, we discovered that training on our entire dataset led to worse performance (higher loss) on most of the higher observability data (LABEL:fig:figure_3_e). Training on only higher observability data was unsurprisingly better for higher observability data, but led to overfitting and worse performance on the low-to-medium observability data (LABEL:fig:figure_3_e). We found that training on between the 40% and 70% most observable data in our dataset led to the best overall performance of the ANN (LABEL:fig:figure_3_e). Altogether, we found that sorting data by observability allowed the curation of a smaller, but richer, dataset that can increase ANN performance (LABEL:fig:figure_3_e). Furthermore, we found that observability was a good predictor of ANN performance and correlated with the ANN error variance (LABEL:fig:figure_3_f). In theory, an ANN should not be able to outperform the minimum error variance $\mathcal{F}^{-1}_{\zeta}$, but in practice an ANN can learn certain heuristics that are viable within the space of the training data, especially for bounded circular data. For instance, when the wind speed is much greater than the ground speed in our flying agent, the wind direction can be approximated as $\zeta\approx\psi+\gamma$, even during unobservable straight flight.
@@
To know when to expect accurate or inaccurate estimates from $\mathcal{H}_{i}$, it is also necessary to know the observability of ${x}_{i}$. However, calculating the true observability $\mathcal{F}^{-1}_{x_{i}}$ using BOUNDS or other empirical observability tools relies heavily on knowing the full state $\mathbf{x}_{k}$ in advance. The agent only has access to $\mathbf{Y}_{\omega_{k}}$, $\mathbf{U}_{\omega_{k}}$, and $\mathcal{R}_{\omega_{k}}$, therefore $\mathbf{x}_{k}$ is not directly available. Even $\mathcal{H}_{i}$’s for each state variable could not be relied on to estimate $\mathbf{x}_{k}$, unless $\mathbf{x}_{k}$ was always observable, thus we have a circular problem that does not allow us to compute observability using traditional approaches. In practice, calculating observability online may be impractical, or impossible. However, we take a data-driven approach to solve this problem by constructing an observability estimator $\mathcal{G}_{i}$ of the form
@@
where $\check{\mathcal{F}}^{-1}_{x_{i,k}}$ is an estimate of the current observability of ${x}_{i}$ at time step $k$. We can similarly use an ANN to construct $\mathcal{G}_{i}$—however, for simple active sensing motifs there may be a relationship between a measured movement, e.g. turning, and observability that can be exploited to develop a less complex model. With sufficiently rich training-data, $\mathcal{G}_{i}$ should produce accurate observability estimates. In some cases, there may be simple observable motifs encoded in $\mathbf{Y}_{\omega_{k}}$ that $\mathcal{G}_{i}$ can easily learn. However, there may be instances where $\mathbf{Y}_{\omega_{k}}$ simply does not contain enough information to precisely estimate ${\check{\mathcal{F}}}^{-1}_{x_{i}}$—therefore, the performance of $\mathcal{G}_{i}$ should be carefully validated.
We trained a wind direction observability estimator ANN (LABEL:fig:figure_supp_ANN_noise_c) using the entire distribution of our training data. For low measurement noise, the estimator was sufficient to estimate the true observability within an order of magnitude and correctly predicted the spikes in observability associated with turns (LABEL:fig:figure_supp_ANN_noise_d). However, larger measurement noise led to overestimates (LABEL:fig:figure_supp_ANN_noise_d). While it may be possible to improve the ANN sensitivity to measurement noise via improved layer design and training data, estimating observability in high-noise data will always be a significant challenge because it is difficult to tease apart “good” changes in measurements due to active sensing and “bad” changes due to noise.
@@
The final step for observability-informed state estimation is to leverage the observability estimate $\check{\mathcal{F}}^{-1}_{x_{i}}$ to adaptively filter the “raw” state estimate $\check{x}_{i,k}$. The filtered state estimate $\hat{x}_{i,k}$ should ideally not change too quickly when observability is low and $\check{x}_{i,k}$ is inaccurate, but converge quickly to $\check{x}_{i,k}$ when observability is high and $\check{x}_{i,k}$ is accurate. The simplest approach is to use an adaptive low-pass filter with coefficients that are dynamically scaled by the estimated observability. We propose a simple update equation,
@@
With validated state and observability estimators (Figure S3), we next applied our observability filter to a testing trajectory with time-varying wind direction, which the ANN’s had not been trained on (LABEL:fig:figure_3_g). This trajectory was designed to only have sporadic bouts of active sensing (turns) to make estimation more challenging. We also injected sensor noise into the measurements $\mathbf{Y}_{\omega_{k}}$. As expected, we found that the ANN raw estimate ($\check{\zeta}$) was only accurate during turns (periods of high observability)—in between turns the estimate was consistently offset (LABEL:fig:figure_3_g). However, the filtered estimate ($\hat{\zeta}$) smoothly and quickly converged to the true value of wind direction ($\zeta$}) during a turn (LABEL:fig:figure_3_g). With sufficiently frequent turns the observability-filtered estimate could track the changing wind direction, although high enough measurement noise did lead to worse performance (LABEL:fig:figure_supp_ANN_noise_eLABEL:fig:figure_supp_ANN_noise_g).
@@
Figure 4: Augmented Information Kalman Filter. a. Overview of the Augmented Information Kalman Filter (AI-KF) state estimation framework. The measurement vector $\mathbf{y}_{k}$ is augmented with a data-driven state estimate $\check{x}_{i,k}$. The corresponding state noise covariance is estimated $\check{\mathcal{R}}_{i,k}$ and used to dynamically adjust the full augmented noise covariance matrix $\mathcal{R}^{\prime}_{k}$ based on the state estimate accuracy (or observability). b. (left) The AI-KF framework is applied to estimate altitude for a simulated trajectory consisting of straight flight with a brief deceleration and sequential acceleration. Altitude is observable during acceleration/deceleration (see LABEL:fig:figure_2_cv) given measurements of forward optic flow, forward acceleration, and vertical acceleration. (right) Noisy measurements used in our simulation with additive zero-mean Gaussian noise with a variance of $10^{-2}$ for each measurement. There is also a brief period of biased (nonzero-mean) noise added to the vertical acceleration measurement. c. A comparison of an Unscented Kalman Filter (UKF, purple) and Augmented Information Unscented Kalman Filter (AI-UKF, green) state estimation results with varied initial altitude state estimates, for the trajectory in b. The shading around the curve insinuates the 2-$\sigma$ estimate bound. The noise variance of the augmented altitude measurement $\mathrm{Var}(\check{z})$ (top, pink scaled) was dynamically adjusted proportionally to the mean acceleration magnitude in the time window $\omega$. d. A comparison of the median error of the UKF and AI-UKF for varying process noise covariance matrix ($Q$) scales and varying initial state covariance matrix ($P_{0}$) scales for varying values of the initial altitude estimate. The results in d correspond to the square with a bold outline. e. Comparison of the UKF and AI-UKF median error across different mean accelerations for varying values of the initial altitude estimate. The AI-UKF always performs better or equal to the vanilla UKF. The advantage of the AI-UKF is most pronounced for small but non-zero levels of acceleration.
@@
While the observability-filtering approach we outline is adequate for estimating individual state variables given sufficient bouts of active sensing, it lacks the powerful statistical rigor of model-based methods like the Kalman filter (KF). For instance, the observability-filter essentially acts like a zero-order hold during periods of low observability, but a KF can potentially make use of “dead-reckoning” to track states with reasonable accuracy in these same periods Brossard2020.
Conversely, Kalman filtering comes with its own issues when dealing with weakly observable and nonlinear systems Butcher2017, Bocquet2017, Devos2021. A major issue is that an incorrect guess for the initial state $\hat{\mathbf{x}}_{k=0}$ can drive the new KF estimate in the incorrect direction. Both the new state prediction and innovation terms in the KF update equation,
$\underbrace{\hat{\mathbf{x}}_{k+1}}_{\begin{subarray}{c}\text{new state}\end{subarray}}=\underbrace{f(\hat{\mathbf{x}}_{k},\mathbf{u}_{k})}_{\begin{subarray}{c}\text{predicted state}\end{subarray}}+\underbrace{K}_{\begin{subarray}{c}\text{Kalman}\\ \text{gain}\end{subarray}}(\underbrace{\tilde{\mathbf{y}}_{k}-h(\hat{\mathbf{x}}_{k},\mathbf{u}_{k})}_{\begin{subarray}{c}\text{innovation}\end{subarray}}),$ (9)
rely on an accurate $\hat{\mathbf{x}}_{k}$, so the KF will struggle to converge. Additionally, the initial state covariance $P_{k=0}$ influences the computation of the Kalman gain $K$, so a wrong guess for $P_{k=0}$ can also lead to unexpected behavior. For linear and observable systems this is generally not a problem, but when observability is low or sparse in time, and dynamics are nonlinear, a KF filter is far from an optimal estimator. Because only one measurement at a time is used to update the state estimate (Equation 9), if $\hat{\mathbf{x}}_{k=0}$ or $P_{k}$ was wrong in the past, then potentially useful prior information is lost. Alternatively, an estimator that is able to consider the entire time-history of measurements may be able to outperform a KF.
We present a method augmenting the traditional Kalman filter framework with information from data-driven estimators, which we call the Augmented Information Kalman Filter (AI-KF) (LABEL:fig:figure_4_a). The AI-KF works by augmenting a KF with extra information from a data-driven state estimator, and then telling the KF when to trust this information more or less with a time-varying noise covariance matrix that is adjusted based on an observability estimate. The main benefit is that the augmented information considers an arbitrarily long time-history of measurements that can make up for a “bad” initialization of the KF. The AI-KF is primarily applicable for nonlinear and partially observable systems that have some state variables with time-varying observability properties.
@@
In principle, the measurement vector can be augmented with any subset of data-driven state variable estimates $\{\check{x}_{i},\dots,\check{x}_{n}\}$. However, for brevity, we show the process for a single state variable.
Next, the original Kalman filter noise covariance $\mathcal{R}$ is augmented with the noise covariance for the data-driven state estimate $\check{x}_{i}$
$\mathcal{R}^{\prime}=$ $\begin{bmatrix}\mathcal{R}&\operatorname{Cov}(\mathbf{y},\tilde{\check{x}}_{i})\\[4.0pt] \operatorname{Cov}(\mathbf{y},\tilde{\check{x}}_{i})^{\top}&\operatorname{Var}(\tilde{\check{x}}_{i})\end{bmatrix}$
where $\tilde{\check{x}}_{i}$ is the data-driven estimation error for the state variable $x_{i}$. The augmented portion of $\mathcal{R}^{\prime}$, which we will refer to as $\check{\mathcal{R}}_{i}$, encodes the covariance of $\check{x}_{i}$ both with itself and all other measurements.
The critical step is constructing $\check{\mathcal{R}}_{i}$ to vary in time based on the accuracy of the state estimate $\check{x}_{i}$. This can be accomplished multiple ways. If $\mathcal{H}_{i}$ is efficient, then the augmented portion of $\mathcal{R}^{\prime}$ is equivalent to the observability from $\mathcal{F}^{-1}$, which can be estimated with $\mathcal{G}_{i}$. However, it is more statistically consistent to construct a covariance function by performing a post hoc analysis of the state estimator $\mathcal{H}_{i}$ itself to determine the error covariance as a function of the estimator inputs $\check{\mathcal{R}}_{i}(Y_{\omega},U_{\omega})$. Notably, the off-diagonal elements of $\check{\mathcal{R}}_{i}$ will generally be non-zero because $\check{x}_{i}$ is a function of $\mathbf{y}_{k}$ (in addition to prior measurements), but the diagonal element $\check{\mathcal{R}}_{i,i}$ will generally have the most impact.
@@
We implemented our AI-KF framework for a flying agent to estimate altitude $z$ and forward velocity $v_{x}$ from only forward ventral optic flow $r_{x}$, forward acceleration $\dot{v}_{x}$, and vertical acceleration $\dot{v}_{z}$ (see Methods). Observability analysis tells us that this estimation problem is solvable given non-zero acceleration (LABEL:fig:figure_2_cv) VanBreugel2014a. However, it is a more challenging problem than wind direction estimation (see difference in observability in LABEL:fig:figure_2_c) and more clearly illustrates the benefit of the AI-KF. In this example, we augment an Unscented Kalman filter (UKF), which is appropriate for nonlinear dynamics—although our method is applicable to most Kalman filter extensions. We show that the AI-UKF can improve state estimation for both simulated data and real data from an outdoor quadcopter flight.

### v2

@@ Results
Adding apparent airflow magnitude ($a$) made offset turns sufficient for wind direction and enabled ground speed estimation for many motifs (LABEL:fig:figure_2_cii). Adding optic flow magnitude ($r$) as well made acceleration and deceleration motifs viable for estimating both wind direction and ground speed, and rendered altitude observable across all motifs (LABEL:fig:figure_2_ciii). Replacing the optic flow measurement with a direct ground speed measurement ($g$) rendered the wind direction and ground speed observable without any motif (even during straight flight) (LABEL:fig:figure_2_civ).
@@
Finally, we imitted the airflow sensor and examined a set consisting of optic flow magnitude ($r$) and direction ($\beta$) and acceleration magnitude ($q$) and direction ($\eta$), which proved suitable for estimating ground speed and altitude during acceleration/deceleration and offset turns (LABEL:fig:figure_2_cv). However, wind direction was never observable, and heading turns had a minimal effect on the observability (LABEL:fig:figure_2_cv).
Table 1: Observable state variables for different sensor sets. For each sensor set from LABEL:fig:figure_2_c, we list the motion primitives (motifs) that can each independently render wind direction ($\zeta$), ground speed ($g$), and altitude ($z$) observable.
Sensor Set Symbols Motif Obs. states
$\zeta$ $g$ $z$
[…]
Although we use a flying agent to provide a concrete example, our results highlight a general principle applicable to many nonlinear partially observabile systems: observability is state-variable-specific and sensor-configuration-dependent. Unlike classical observability analyses designed to reveal the combinations of states that are most observable (by analyzing the eigenvectors of the observability Gramian $\mathcal{O}^{T}\mathcal{O}$), BOUNDS reveals how optimal active sensing strategies can be uniquely tailored to target specific state variables. The motifs BOUNDS reveals can subsequently be used either to generate hypotheses about the role motifs may serve in organisms, or to inform control/estimation strategies and algorithms for autonomous agents with limited sensors. The same approach we show in our case study can be applied in a similar to any arbitrary system to discover measurement- and movement-dependent sensing motifs.
@@
Figure 3: Observability-informed state estimation. a. Artificial neural network (ANN) state estimator for wind direction ($\mathcal{H}_{\zeta}$). Inputs are the yaw heading angle ($\psi$), apparent wind angle ($\gamma$), and course direction angle ($\beta$) in the time window $\omega$. b. A subset of simulated flying agent trajectories used for training and testing the ANN state estimator. $N=40,000$ trajectories total. c. The distribution (bottom) and cumulative distribution function (top) of wind direction observability ($\mathcal{F}^{-1}_{\zeta}$) across all training and testing trajectories. d. A comparison of the testing error variance for the wind direction state estimator ANN ($\mathcal{H}_{\zeta}$) when trained on different percentages of the training data—for data sorted by observability (sorted high-to-low, top) and randomly sorted (bottom). The y-axis indicates how much data was used in training, and the x-axis indicates what bin percentile was used to test the ANN (testing data always sorted by observability). e. Same as d. but showing the relative change in error variance between observability-sorted and randomly sorted training data. Red values (<< 0) indicate better performance when sorting by observability. f. The mean observability $\text{mean}(\mathcal{F}^{-1}_{\zeta})$ of each bin in the training dataset vs the ANN error variance $\text{Var}(\check{\zeta}-\zeta)$ for ANNs trained on different percentages of the training dataset for observability-sorted (left) and randomly sorted (right) training data. The theoretical bound of the ANN performance is shown as a dashed line (bounded at 1 for circular data; see Methods). g. (top) An example simulated trajectory of a flying agent in the presence of time-varying wind direction. Color along trajectory indicates observability. (bottom) A comparison of true wind direction ($\zeta$}), ANN wind direction estimate ($\check{\zeta}$), and the observability-filtered estimate ($\hat{\zeta}$). Measurement noise covariance was set to $\mathcal{R}=10^{-3}I_{p\times p}$.
@@
How do we exploit the sporadic information provided by active sensing maneuvers? Here we develop a simple framework that leverages knowledge of the observability to determine when state estimates are likely to be accurate, and when they are not, and use this knowledge to generate an observability filtered estimate. Building such a filter involves four steps: constructing a state estimator, curating data by observability, training an observability estimator, and observability-informed filtering. We show an example application of each step for wind direction estimation.
@@
$\mathcal{H}_{i}$ is a function that maps the time-history of measurements $\mathbf{Y}_{\omega_{k}}$ (and optionally inputs $\mathbf{U}_{\omega_{k}}$ and measurement noise covariance $\mathcal{R}_{\omega_{k}}$) in the time window $\omega$ to the current state estimate $\check{x}_{i,k}$ at time step $k$. Notably, $\mathcal{H}_{i}$ can consider the entire time window $\omega$ at once, and is not sensitive to initialization, thus can potentially avoid the pitfalls associated with Kalman filtering. In principle, $\mathcal{H}_{i}$ could be any function, but we employ a feed-forward artificial neural network (ANN) for its generality, flexibility, and ability to operate without relying on a predefined dynamical model. We designed an ANN for estimating wind direction using angular measurements $\psi,\gamma,\beta]$ (LABEL:fig:figure_3_a, LABEL:fig:figure_supp_ANN_noise_a), which yields accurate estimates LABEL:fig:figure_supp_ANN_noise_b, but only during heading changes as predicted by our observability analysis (LABEL:fig:figure_2_ci).
@@
The distribution of observability in the data used to train the ANN is critical: training on primarily unobservable data will yield poor results. Knowledge of active sensing motifs discovered by BOUNDS can help us selectively generate high quality training data, or curate a subset of an existing dataset, to efficiently train an accurate estimator.
To train our wind direction estimator, we randomly simulated a diversity of trajectories spanning a large envelope of the state space and with a wide distribution of observability values (LABEL:fig:figure_3_bLABEL:fig:figure_3_c, see Methods). We separated the data into ten equally sized bins, sorted by the observability of wind direction (LABEL:fig:figure_3_c). This allowed us to train ANNs on inputs with distinct distributions of observability. We analyzed the ANN wind direction estimate error variance $\text{Var}(\zeta-\check{\zeta})$ when trained on different percentages of the training dataset—once when the training data was sorted by observability level, and once when the data was randomly sorted (LABEL:fig:figure_3_d, Figure S4).
Sorting data by observability allowed the curation of a smaller, but richer, dataset that increased ANN performance. Training on between the 40% and 70% most observable data led to the best overall performance (LABEL:fig:figure_3_e). Furthermore, observability was a good predictor of ANN performance and correlated with the ANN error variance (LABEL:fig:figure_3_f), though our ANN outperformed the theoretical bound. This is likely a result of the ANN learning biased heuristics from the training data. For instance, when the wind speed is much greater than the ground speed in our flying agent, the wind direction can be approximated as $\zeta\approx\psi+\gamma$, even during unobservable straight flight.
@@
Knowing when $\mathcal{H}_{i}$ produces accurate estimates requires knowing the observability of ${x}_{i}$. However, calculating the true observability $\mathcal{F}^{-1}_{x_{i}}$ relies on knowledge of the full state $\mathbf{x}_{k}$ in advance. The agent only has access to $\mathbf{Y}_{\omega_{k}}$, $\mathbf{U}_{\omega_{k}}$, and $\mathcal{R}_{\omega_{k}}$, therefore $\mathbf{x}_{k}$ is not directly available. We take a data-driven approach to solve this problem by constructing an observability estimator $\mathcal{G}_{i}$ of the form
@@
where $\check{\mathcal{F}}^{-1}_{x_{i,k}}$ is an estimate of the current observability of ${x}_{i}$ at time step $k$. We can similarly use an ANN to construct $\mathcal{G}_{i}$—however, for simple active sensing motifs there may be a heuristic relationship between a measured movement, e.g. turning, and observability that can be exploited to develop a less complex model.
We trained a wind direction observability estimator ANN (LABEL:fig:figure_supp_ANN_noise_c) using the entire distribution of our training data. At low noise, the estimator predicted true observability within an order of magnitude, capturing turn-associated spikes (LABEL:fig:figure_supp_ANN_noise_d); higher noise caused overestimates because it is difficult to tease apart “good” changes in measurements due to active sensing and “bad” changes due to noise.
@@
The final step leverages observability estimates $\check{\mathcal{F}}^{-1}_{x_{i}}$ to adaptively filter raw state estimates $\check{x}_{i,k}$. The filtered state estimate $\hat{x}_{i,k}$ should not change quickly when observability is low and $\check{x}_{i,k}$ is inaccurate, but should quickly converge to $\check{x}_{i,k}$ when observability is high and $\check{x}_{i,k}$ is accurate. A simple solution is an adaptive low-pass filter with coefficients dynamically scaled by the estimated observability,
@@
We applied our observability filter to a trajectory with time-varying wind direction, which the ANNs had not been trained on (LABEL:fig:figure_3_g). This trajectory was designed to only have sporadic bouts of active sensing (turns) to make estimation more challenging. We also injected sensor noise into the measurements $\mathbf{Y}_{\omega_{k}}$. As expected, the ANN raw estimate ($\check{\zeta}$) was accurate only during turns; between turns, estimates were consistently offset (LABEL:fig:figure_3_g). However, the filtered estimate ($\hat{\zeta}$) quickly converged to the true value of wind direction ($\zeta$}) during a turn (LABEL:fig:figure_3_g). With sufficiently frequent turns the observability-filtered estimate could track the changing wind direction, although high enough measurement noise did lead to worse performance (LABEL:fig:figure_supp_ANN_noise_eLABEL:fig:figure_supp_ANN_noise_g).
@@
[…]
The observability-filtering approach does not take full advantage of the dynamic relationships between states and inputs, and lacks the statistical rigor of model-based methods like the Kalman filter (KF). The observability-filter essentially acts like a zero-order hold during periods of low observability, but a KF can potentially make use of “dead-reckoning” to track states with reasonable accuracy in these same periods 10. Kalman filtering, however, comes with its own challenges for weakly observable and nonlinear systems 11, 5: an incorrect guess for the initial state $\hat{\mathbf{x}}_{k=0}$ can drive the new estimate in the wrong direction. Both the predicted state and innovation terms in the KF update equation,
$\underbrace{\hat{\mathbf{x}}_{k+1}}_{\begin{subarray}{c}\text{new state}\end{subarray}}=\underbrace{\mathbf{f}(\hat{\mathbf{x}}_{k},\mathbf{u}_{k})}_{\begin{subarray}{c}\text{predicted state}\end{subarray}}+\underbrace{K}_{\begin{subarray}{c}\text{Kalman}\\ \text{gain}\end{subarray}}(\underbrace{\tilde{\mathbf{y}}_{k}-\mathbf{h}(\hat{\mathbf{x}}_{k},\mathbf{u}_{k})}_{\begin{subarray}{c}\text{innovation}\end{subarray}}),$ (4)
rely on an accurate $\hat{\mathbf{x}}_{k}$, so the KF will struggle to converge when $\hat{\mathbf{x}}_{k}$ is incorrect . Additionally, the initial state covariance $P_{k=0}$ influences the computation of the Kalman gain $K$, so a wrong guess for $P_{k=0}$ can also lead to unexpected behavior. For linear and observable systems this is generally not a problem, but when observability is low or sparse in time, and dynamics are nonlinear, a KF filter is far from an optimal estimator. Because only one measurement at a time is used to update the state estimate (Equation 4), if $\hat{\mathbf{x}}_{k=0}$ or $P_{k}$ was wrong in the past, then potentially useful prior information is lost. Alternatively, an estimator that is able to consider the entire time-history of measurements—such as our $\mathcal{H}$ function—may be able to outperform a KF.
Our Augmented Information Kalman Filter (AI-KF) augments traditional Kalman filtering with data-driven state estimates from $\mathcal{H}$ (LABEL:fig:figure_4_a). The AI-KF works by augmenting a KF with extra information from a data-driven state estimator, and then telling the KF when to trust this information more or less with a time-varying noise covariance matrix that is adjusted based on an observability estimate. Unlike heuristic ANN-KF hybrids 34, 18, 25, 32, the AI-KF weights augmented measurements by their Cramér-Rao bound, grounding information fusion in observability theory. The main benefit is that the augmented information considers an arbitrarily long time-history of measurements that can make up for a “bad” initialization of the KF. This observability-based weighting also enables interpretability, making when and why the AI-KF outperforms traditional filtering explicit.
@@
In principle, the measurement vector can be augmented with any subset of data-driven state variable estimates $\check{x}_{0},\dots,\check{x}_{n}$}. However, for brevity, we show the process for a single state variable. Next, the original Kalman filter noise covariance $\mathcal{R}$ is augmented with the noise covariance for the data-driven state estimate $\check{x}_{i}$
$\mathcal{R}^{\prime}=$ $\begin{bmatrix}\mathcal{R}&\operatorname{Cov}(\mathbf{y},\tilde{\check{x}}_{i})\\[4.0pt] \operatorname{Cov}(\mathbf{y},\tilde{\check{x}}_{i})^{\top}&\operatorname{Var}(\tilde{\check{x}}_{i})\end{bmatrix}$,
where $\tilde{\check{x}}_{i}$ is the data-driven estimation error for the state variable $x_{i}$. The augmented portion of $\mathcal{R}^{\prime}$, which we will refer to as $\check{\mathcal{R}}_{i}$, encodes the covariance of $\check{x}_{i}$ both with itself $\big(\operatorname{Var}(\tilde{\check{x}}_{i})\big)$ and all other measurements $\big(\operatorname{Cov}(\mathbf{y},\tilde{\check{x}}_{i})\big)$.
The critical step is constructing $\check{\mathcal{R}}_{i}$ to vary in time based on the accuracy of the state estimate $\check{x}_{i}$. This task can be accomplished multiple ways. If $\mathcal{H}_{i}$ is unbiased and efficient [^2] An efficient (unbiased) estimator is an estimator for which the error variance reaches the Cramér–Rao Bound20., then the augmented portion of $\mathcal{R}^{\prime}$ is equivalent to the observability from $\mathcal{F}^{-1}$, which can be estimated with $\mathcal{G}_{i}$. However, it is more statistically consistent to construct a covariance function by performing a post hoc analysis of the state estimator $\mathcal{H}_{i}$ itself to determine the error covariance as a function of the estimator inputs $\check{\mathcal{R}}_{i}(Y_{\omega},U_{\omega})$. Notably, the off-diagonal elements of $\check{\mathcal{R}}_{i}$ will generally be non-zero because $\check{x}_{i}$ is a function of $\mathbf{y}_{k}$ (in addition to prior measurements), but the diagonal element $\check{\mathcal{R}}_{i,i}$ will have the most impact. Because the Kalman filter assumes that each measurement (including $\check{x}_{i}$ is independent of the prior state estimate (including $\hat{x}_{i-1}$), there is a potential for double counting information once the KF has converged to an accurate state estimate. Although this will not harm the accuracy of the state estimate, it can artificially decrease the error variance. In the methods we describe a simple resolution to this that allows the AI-KF to converge to a classic KF when the two are in agreement.
@@
We implemented our AI-KF framework for a flying agent to estimate altitude $z$ and forward velocity $v_{x}$ from only forward ventral optic flow $r_{x}$, forward acceleration $\dot{v}_{x}$, and vertical acceleration $\dot{v}_{z}$ (see Methods). Observability analysis tells us that this estimation problem is solvable given non-zero acceleration (LABEL:fig:figure_2_cv) 59. It is a more challenging problem than wind direction estimation (see difference in observability in LABEL:fig:figure_2_c) and more clearly illustrates the benefit of the AI-KF. In this example, we augment an Unscented Kalman filter (UKF), which is appropriate for nonlinear dynamics—although our method is applicable to most Kalman filter extensions.